# Model Merging in LLMs, MLLMs, and Beyond: Methods, Theories, Applications and Opportunities

### v4

@@ Abstract
Model merging is an efficient empowerment technique in the machine learning community that does not require the collection of raw training data and does not require expensive computation. As model merging becomes increasingly prevalent across various fields, it is crucial to understand the available model merging techniques comprehensively. However, there is a significant gap in the literature regarding a systematic and thorough review of these techniques. This survey provides a comprehensive overview of model merging methods and theories, their applications in various domains and settings, and future research directions. Specifically, we first propose a new taxonomic approach that exhaustively discusses existing model merging methods. Secondly, we discuss the application of model merging techniques in large language models, multimodal large language models, and 10+ machine learning subfields, including continual learning, multi-task learning, few-shot learning, etc. Finally, we highlight the remaining challenges of model merging and discuss future research directions. A comprehensive list of papers about model merging is available at https://github.com/EnnengYang/Awesome-Model-Merging-Methods-Theories-Applications.
1 Introduction
Model merging, also known as model fusion, is an effective technique that merges the parameters of multiple separate models with different capabilities to build a universal model without needing access to the original training data or expensive computation. The concept most relevant to model merging is ensemble learning [36, 159, 205, 122], as both facilitate knowledge fusion and transfer. As shown in Figure 1, the main difference between them is that ensemble learning must save all individual models and fuse the predictions (or outputs) of multiple models during the inference phase, whereas model merging performs merging directly at the parameter level and only has one final model during inference. This gives model merging more attractive properties.
@@ 1 Introduction
Although model merging is a relatively young topic, it is evolving rapidly and has already found applications in several domains. For example, in foundation models, models fine-tuned by different downstream tasks are merged to enhance the capabilities of large language models, and image generative models with different styles are merged to create a new model with mixed-style capabilities. In particular, the number of pre-trained and fine-tuned checkpoints in the machine learning community has grown exponentially in recent years, including open-source repositories such as Huggingface [207], torchvision [124], and timm [206], making it easy for users to obtain well-trained expert models of varying abilities. These rich model repositories further promote the rapid development of model merging direction.
Figure 1: An illustration of the ensemble learning paradigm versus the model merging paradigm. (a) $T$ separate models for $T$ tasks, (b) Assemble $T$ separate models for $T$ tasks, (c) A merged model for $T$ tasks.
@@
As model merging becomes increasingly popular in various areas of the machine learning community, it is crucial to have a comprehensive understanding of the advantages and limitations of existing model merging techniques and their applications across different domains. Although some efforts have been made by the community [108, 239, 177, 52], there are still large gaps to be filled. More specifically, MergeKit [52] and FusionBench [177] are technical reports in which only seven representative methods are discussed in MergeKit, and eight merging methods are discussed in FusionBench. Additionally, Zheng et al. [239] discuss the topic of “learning from models” and it only mentions model merging as a subsection (single page only) in the whole paper. The most related work to the “model merging” topic is [108], but in terms of application, it only discusses model merging in three scenarios: federated learning, fine-tuning, and distillation. It also ignores a lot of recently published articles due to the rapid evolution of the model merging direction. To address these gaps, this survey aims to elucidate the methods, theories, applications, and future trends in model merging direction, providing a comprehensive classification of relevant approaches. In particular, this paper enhances the comprehensive understanding of model merging by covering three main aspects:
@@
First, how are existing model merging methods classified? We first propose a new taxonomy in Figure 2 (upper part) that divides existing model merging methods into two phases (§2): pre-merging and during-merging. (i) Pre-merging methods aim to create better conditions for merging. It is further divided into using linearized fine-tuning to achieve weight space and input space disentanglement, performing architectural transformations to convert heterogeneous models into homogeneous models, and aligning weights to place them in the same basin. (ii) During-merging methods focus on designing sophisticated techniques to merge multiple models into one. These methods address task conflict and interference problems when merging models. They can be further divided into basic merging methods that perform the simplest parameter merging strategy; weighted merging methods that merge multiple models according to the importance calculated by specific rules; subspace merging methods that project multiple models into sparse subspaces for merging; routing-based methods that dynamically merge models according to input samples during inference; and the post-calibration based method that corrects the merged model. In addition to these methods, we also discuss the theoretical or empirical analysis of model merging.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ Model Merging: Methods, Theories, Applications , ver, color=hidden-draw, fill=mygray!100, text width=19.5em, text=black [ Methods (§2), fill=myred!80, text width=8.0em, text=black [ Pre-Merging Mehtods (§2.2), fill=myred!60, text width=14em, text=black [ Linearization Fine-tuning (§2.2.1)
Architecture Transformation (§2.2.2)
Weight Alignment (§2.2.3) , color=hidden-draw, fill=myred!40, text width=23em, text=black ] ] [ During-Merging Methods (§2.3), fill=myred!60, text width=14em, text=black [ Basic Merging Methods (§2.3.1)
Weighted-based Merging Methods (§2.3.2)
Subspace-based Merging Methods (§2.3.3)
Routing-based Merging Methods (§2.3.4)
Post-calibration based Methods (§2.3.5) , color=hidden-draw, fill=myred!40, text width=23em, text=black ] ] [ Theories and Analysis (§2.4), fill=myred!60, text width=14em, text=black ] ] [ Applications (§34), fill=myblue!80, text width=8.0em, text=black [ Large Language Models (§3.1), , fill=myblue!60, text width=17em, text=black [ Human Value Alignment for LLMs (§3.1.1)
Detoxifcation of LLMs (§3.1.2)
Knowledge Unlearning of LLMs (§3.1.3)
Faster Training of LLMs (§3.1.4)
Combine the Capabilities of Expert LLMs (§3.1.5) , color=hidden-draw, fill=myblue!40, text width=20em, text=black ] ] [ Multimodal Large Language Models (§3.2), , fill=myblue!60, text width=17em, text=black [ Multimodal Fusion (§3.2.1);
Cross-modal Knowledge Transfer (§3.2.2) , color=hidden-draw, fill=myblue!40, text width=20em, text=black ] ] [ Visual Generative Models (§3.3), , fill=myblue!60, text width=17em, text=black [ Style Mixing (§3.3.1)
Reducing Training Cost (§3.3.2)
Enhancing the Faithfulness (§3.3.3) , color=hidden-draw, fill=myblue!40, text width=20em, text=black ] ] [ Continual Learning (§4.1), fill=myyellow!60, text width=15em, text=black [ Mitigate Catastrophic Forgetting (§4.1.1) , color=hidden-draw, fill=myyellow!40, text width=22em, text=black ] ] [ Multi-Task/Domain/Objective/Auxiliary Learning (§4.2), fill=myyellow!60, text width=22em, text=black [ Knowledge Transfer in MTL (§4.2.1)
Knowledge Transfer in MOO (§4.2.2)
Knowledge Transfer in MDL (§4.2.3)
Knowledge Transfer in ATL (§4.2.4) , color=hidden-draw, fill=myyellow!40, text width=15em, text=black ] ] [ Out-of-Distribution/Domain Generalization (§4.3), fill=myyellow!60, text width=20.5em, text=black [ Better OOD Generalization (§4.3.1)
Better DG Generalization (§4.3.2) , color=hidden-draw, fill=myyellow!40, text width=16.5em, text=black ] ] [ Federated Learning (§4.4), fill=myyellow!60, text width=15em, text=black [ Local Knowledge Aggregation (§4.4.2) , color=hidden-draw, fill=myyellow!40, text width=22em, text=black ] ] [ Zero-shot/Few-Shot Learning (§4.5), fill=myyellow!60, text width=15em, text=black [ Zero-shot Knowledge Transfer (§4.5.1)
Few-shot Knowledge Transfer (§4.5.2) , color=hidden-draw, fill=myyellow!40, text width=22em, text=black ] ] [ Adversarial Learning (§4.6), fill=myyellow!60, text width=15em, text=black [ Model Attack (§4.6.1)
Model Defense and Copyright Protection (§4.6.2) , color=hidden-draw, fill=myyellow!40, text width=22em, text=black ] ] ] ]
@@
Figure 2: The taxonomy of model merging in machine learning. This general framework covers advanced model merging methods and theories (top part), as well as practical applications of model merging techniques to foundation models and more than 10 machine learning subfields (bottom part).
@@
Second, which applications can benefit from model merging? We discuss in detail the various use cases of model merging in foundation models (§3) and over ten subfields of machine learning (§4). As shown in Figure 2 (lower part), model merging can be applied to a variety of foundation models, including large language models, multimodal large language models, and visual generative models. For example, model merging in large language models can help mitigate untruthfulness and toxicity output, accomplish knowledge unlearning, and speed up training. Moreover, model merging also arises in different machine learning subfields, such as continual learning, multi-task/multi-domain learning, few-shot learning, and other subfields, to solve a variety of challenges. For instance, in continual learning, model merging can mitigate catastrophic forgetting of old tasks. In multi-task learning, multi-objective learning and multi-domain learning, it facilitates knowledge transfer. Additionally, in adversarial learning, model merging can be employed for both attack and defense strategies.
@@
  • Methodology Overview: We provide a comprehensive summary of the technical aspects of model merging. Specifically, we propose a new taxonomy that divides existing model merging methods into two stages and further subdivides the methods in each stage according to key techniques. Additionally, we discuss theoretical analysis work related to model merging.
  • Application Overview: We offer a comprehensive summary of the application aspects of model merging. Specifically, we explore the application of model merging to foundation models and $10+$ machine learning subfields, demonstrating how model merging can address existing challenges in these areas.
  • Future Directions: We outline several remaining challenges and future directions for model merging. We believe that model merging needs to be further explored in the future from the perspectives of performance gap, theoretical analysis, trustworthy guarantees, cross-disciplinary applications, etc.
The main structure of this paper is as follows: §1 is an introduction, and §2 offers a comprehensive discussion of advanced model merging methods from a technical perspective. In §3 and §4, we summarize the applications of model merging in various foundation models and different subfields within machine learning, respectively. The remaining challenges and future research directions are discussed in §5. Finally, we conclude this paper in §6.
2 Advanced Model Merging Methods
In this section, we first introduce the notation and problem definition of model merging in §2.1. We then elaborate on advanced model merging methods (Table 1 summarizes the primary purpose of each category of methods). Existing model merging techniques can be roughly divided into the following two categories: (i) Before Merging Methods in §2.2: it provides better prior knowledge for model merging. (ii) During Merging Methods in §2.3: it resolves task conflict/interference by various strategies, and then performs parameter merging operations. Finally, we conclude with theories or explanations for the effectiveness of model merging in §2.4.
Table 1: A summary of existing model merging methods.
@@ 2 Advanced Model Merging Methods
Linearization Fine-tuning (§2.2.1) Disentangling different models in input space and weight space
@@
Weight Alignment (§2.2.3) Repermutate multiple models into the same basin
@@
Subspace-based Merging Methods (§2.3.3) Merge multiple models by projecting them into a sparse subspace
@@
2.1 Notation and Model Merging Problem Definition
@@
In this survey, we focus on parameter-wise merging. In other words, the goal of model merging is to merge the parameters $\small\{\Theta^{(1)},\ldots,\Theta^{(T)}\}$, and finally obtain the new parameters $\small\Theta^{(merge)}=\texttt{merge}(\Theta^{(1)},\ldots,\Theta^{(T)})$. One straightforward solution for merging models is weighted averaging [190, 164], defined as $\small\Theta^{(merge)}=\frac{1}{T}\sum_{t=1}^{T}\Theta^{(t)}$. However, the performance of this approach is often unacceptably poor or infeasible due to several possible factors: (i) The lack of suitable merging conditions, such as multiple models not being in the same basin or having inconsistent architectures. (ii) There are conflicts and interference among multiple models. We illustrate how advanced methods address these issues in §2.2 and §2.3, respectively.
2.2 Pre-Merging Methods
To provide better preconditions for model merging, one class of work focuses on the fine-tuning step of independent models, such as fine-tuning the linearized model instead of the nonlinear model (in §2.2.1). Additionally, when multiple model architectures that need to be merged are inconsistent, they must be pre-transformed to the same architecture (in §2.2.2). Finally, another class of work attempts to align the weights/parameters before merging (in §2.2.3).
2.2.1 Linearization Fine-tuning
Ortiz-Jimenez et al. [136] reveal that one necessary condition for effective model merging is ‘weight disentanglement’. This means that different directions of the weight space correspond to functional changes in disjoint regions of the input space. For example, if model $1$ in weight space corresponds to a function change on $\small\mathcal{D}_{1}$ in input space and model 2 in weight space corresponds to a function change on $\small\mathcal{D}_{2}$ in input space, then the merged model will not interfere with each other in terms of function change. In other words, two models satisfying this weight disentanglement property can coexist in one model without affecting their respective performance, which is a very attractive property.
@@ 2.2.1 Linearization Fine-tuning
To achieve weight disentanglement, Ortiz-Jimenez et al. [136] propose fine-tuning the linearized model along the tangent space [74] of the pre-trained model during the fine-tuning stage, rather than in the original space of the nonlinear model. However, linearized fine-tuning with all parameters is more expensive than nonlinear fine-tuning. To accelerate this process, some works suggest linearizing only part of the layers. For example, Tang et al. [180] propose partially linearizing the Adapter modules and then merging Adapters. Jin et al. [85] suggest linearly fine-tuning only the linear layers in the attention modules of the full model. Furthermore, TAFT [118] develops an efficient linearization method for the Transformer [191] architectures, which directly derives closed-form linearized solutions for transformer networks. In summary, fine-tuning in the tangent space makes it easier to disentanglement the input space and weight space, thereby reducing the interference in subsequent model merging.
2.2.2 Architecture Transformation
In some cases, models that need to be merged may have different architectures and cannot be merged directly. To solve this problem, some studies [10, 193, 194, 134] propose to perform architecture transformation before merging, that is, transform multiple models with different architectures into the same architecture as shown in Figure 3 (a). For example, GAN Cocktail [10] attempts to merge multiple GAN models $\{\Theta^{(1)},\ldots,\Theta^{(T)}\}$ with different architectures. It transforms all GAN models $\small\Theta^{(t)}$ ($t\in\{1,2,\ldots,T\}\;\&\;t\!\neq\!k$) into a specified target model $\Theta^{(k)}$, that is, $\small\Theta^{(k)}$ is used as the initialization to learn the output of $\small\Theta^{(t)}$, while adding implicit regularizations to ensure that $\small\Theta^{(k)}$ does not forget knowledge of the task $k$. Consequently, the transformed GAN models have the same structure and shared knowledge, facilitating further model merging. Similarly, FuseChat [194] proposes to merge chat LLMs with diverse architectures and scales (e.g., NH2-Mixtral-8x7B [83], NH2-Solar-10.7B [93], OpenChat-3.5-7B [195] in their practical applications). Specifically, FuseChat first uses knowledge distillation to transform all the architectures to match that of OpenChat-3.5-7B, and then performs the model merge operation. Unlike the above distillation-based approach, CLAFusion [134] adds layers/blocks (with weights set to the identity matrix) to the smaller model to align its architecture with that of the larger model. In summary, merging models with different architectures requires first transforming all models into a common architecture to merge later.
@@ 2.2.2 Architecture Transformation
Figure 3: (a) An illustration of an architectural transformation that transforms multiple heterogeneous models into homogeneous models, allowing subsequent direct parameter-level merge operations to be performed. (b) An illustration of the weights/parameters alignment, that is, permuting the neural network model $\small\Theta^{(1)}$ so that it aligns with the model $\small\Theta^{(2)}$.
2.2.3 Weight Alignment

### v5

@@ Abstract.
Model merging is an efficient empowerment technique in the machine learning community that does not require the collection of raw training data and does not require expensive computation. As model merging becomes increasingly prevalent across various fields, it is crucial to understand the available model merging techniques comprehensively. However, there is a significant gap in the literature regarding a systematic and thorough review of these techniques. This survey provides a comprehensive overview of model merging methods and theories, their applications in various domains and settings, and future research directions. Specifically, we first propose a new taxonomic approach that exhaustively discusses existing model merging methods. Secondly, we discuss the application of model merging techniques in large language models, multimodal large language models, and more than ten machine learning subfields, including continual learning, multi-task learning, few-shot learning, etc. Finally, we highlight the remaining challenges of model merging and discuss future research directions. A comprehensive list of papers about model merging is available at https://github.com/EnnengYang/Awesome-Model-Merging-Methods-Theories-Applications.
Model Merging, Large Language Model, Multimodal Large Language Models, Continual Learning, Multitask Learning
copyright: acmlicensedjournalyear: 0000doi: 0000000.0000000journal: JACMjournalvolume: 00journalnumber: 0article: 000publicationmonth: 0ccs: Computing methodologies Machine learningccs: Computing methodologies Learning paradigms
1. Introduction
Model merging, also known as model fusion, is an effective technique that merges the parameters of multiple expert models with different capabilities to build a universal model without needing access to the original training data or expensive computation. The concepts most relevant to model merging are multi-task learning (Caruana, 1997; Vandenhende et al., 2021), federated learning (Li et al., 2020), and ensemble learning (Dietterich et al., 2002; Sagi and Rokach, 2018), as all aim to facilitate knowledge fusion and transfer. Specifically, as shown in Figure 1, single-task learning trains a separate expert model for each task; while this avoids interference across tasks, it hinders knowledge transfer and incurs high storage costs. Multi-task learning maintains a single shared model, but requires collecting data from all tasks to jointly update the model parameters. Federated learning trains local models on clients’ private data and then aggregates them on a central server to obtain a global model, but this requires multiple rounds of communication between the server and clients. Ensemble learning aggregates the outputs of multiple expert models; however, it must store all expert models, leading again to substantial storage overhead. In contrast, model merging directly aggregates the parameters of several well-trained expert models to fuse their knowledge, making it highly efficient and inexpensive. These properties render model merging a particularly attractive paradigm for knowledge integration.
@@ 1. Introduction
Although model merging is a relatively young topic, it is evolving rapidly and has already found applications in several domains. For example, in foundation models, models fine-tuned by different downstream tasks are merged to enhance the capabilities of large language models, and visual generative models with different styles are merged to create a new model with mixed-style capabilities. In particular, the number of pre-trained and fine-tuned checkpoints in the machine learning community has grown exponentially in recent years, including open-source repositories such as Huggingface (Wolf et al., 2019), torchvision (Marcel and Rodriguez, 2010), and timm (Wightman, 2019), making it easy for users to obtain well-trained expert models of varying abilities. These rich model repositories further promote the rapid development of the model merging direction.
Figure 1. Schematic comparison of five machine learning paradigms (adapted from (Jin et al., 2023)): (a) Single Task Learning, (b) Multi-Task Learning, (c) Federated Learning, (d) Ensemble Learning, (e) Model Merging. Key attributes are highlighted in red and green.
@@
As model merging becomes increasingly popular in various areas of the machine learning community, it is crucial to have a comprehensive understanding of the advantages and limitations of existing model merging techniques and their applications across different domains and settings. Although some efforts have been made by the community (Li et al., 2023c; Zheng et al., 2023; Tang et al., 2025a; Goddard et al., 2024a), there are still large gaps to be filled. More specifically, MergeKit (Goddard et al., 2024a) and FusionBench (Tang et al., 2025a) are benchmark works in which only seven representative methods are discussed in MergeKit, and eight merging methods are discussed in FusionBench. Additionally, Zheng et al. (2023) discuss the topic of ”learning from models” and it only mentions model merging as a subsection (single page only) in the whole paper. The most related work to the ”model merging” topic is (Li et al., 2023c), but in terms of application, it only discusses model merging in three scenarios: federated learning, fine-tuning, and distillation. It also ignores a lot of recently published articles due to the rapid evolution of the model merging direction. To address these gaps, this survey aims to elucidate the methods, theories, applications, and future trends in model merging direction, providing a comprehensive classification of relevant approaches. In particular, this paper enhances the comprehensive understanding of model merging by covering three main aspects:
@@
First, how are existing model merging methods classified? We first propose a new taxonomy in Figure 2 (upper part) that divides existing model merging methods into two phases (§2): pre-merging and during-merging. (i) Pre-merging methods aim to create better conditions for merging. It is further divided into using linearized fine-tuning to achieve weight space and input space disentanglement or subspace fine-tuning schemes that reduce parameter conflicts, performing architectural transformations to convert heterogeneous models into homogeneous models, and aligning weights to place them in the same basin. (ii) During-merging methods focus on designing sophisticated techniques to merge multiple models into one. These methods address task conflict and interference problems when merging models. They can be further divided into basic merging methods that perform the simplest parameter merging strategy; weighted merging methods that merge multiple models according to the importance calculated by specific rules; subspace merging methods that project multiple models into sparse or low-rank subspaces for merging; optimization-based methods leverage model-specific properties to define optimization objectives that minimize information loss; routing-based methods that dynamically merge models according to input samples during inference; and the post-calibration-based method that corrects the merged model. In addition to these methods, we also discuss the theoretical or empirical analysis of model merging (§3).
[…]
@@
Figure 2. The taxonomy of model merging in machine learning. This general framework covers advanced model merging methods (top part), theoretical and empirical analysis (middle part), and practical applications of model merging techniques to foundation models and more than 10 machine learning subfields (bottom part).
@@
Second, which applications can benefit from model merging? We discuss in detail the various use cases of model merging in foundation models (§4) and over ten subfields of machine learning (§5). As shown in Figure 2 (lower part), model merging can be applied to a variety of foundation models, including large language models, multimodal large language models, and visual generative models. For example, model merging in large language models can help mitigate untruthfulness and toxicity output, accomplish knowledge unlearning, and speed up training. Moreover, model merging also arises in different machine learning subfields, such as continual learning, multi-task learning, few-shot learning, and other subfields, to solve a variety of challenges. For instance, in continual learning, model merging can mitigate catastrophic forgetting of old tasks. In multi-task/objective/domain learning, it facilitates knowledge transfer. Additionally, in adversarial learning, model merging can be employed for both attack and defense strategies.
@@
  • Methodology Overview: We present a taxonomy of model merging methods that splits them into two stages and further classifies each stage by key techniques. We also review existing theoretical analyses and empirical evaluations.
  • Application Overview: We survey applications of model merging in foundation models and more than ten machine learning subfields, showing how it helps address existing challenges.
  • Future Directions: We discuss open challenges and future directions for model merging, including performance gaps, theoretical understanding, trustworthy guarantees, and cross-disciplinary applications.
The structure of this paper is as follows: §1 is an introduction, and §2 offers a comprehensive discussion of advanced model merging methods from a technical perspective. §3 validates representative model merging methods from both theoretical and experimental perspectives. In §4 and §5, we summarize the applications of model merging in various foundation models and different subfields within machine learning, respectively. The remaining challenges and future research directions are discussed in §6. Finally, §7 concludes the paper.
2. Advanced Model Merging Methods
In this section, we first introduce the notation and problem definition of model merging in §2.1. We then elaborate on advanced model merging methods (Table 1 summarizes the primary purpose of each category of methods). Existing model merging techniques can be roughly divided into the following two categories: (i) Before Merging Methods in §2.2: it provides better prior knowledge for model merging. (ii) During Merging Methods in §2.3: it resolves task conflict/interference by various strategies, and then performs parameter merging operations.
Table 1. A summary of existing model merging methods.
@@ 2. Advanced Model Merging Methods
Merger-Friendly Fine-tuning2.2.1) Design fine-tuning strategies that facilitate downstream model merging
@@
Weight Re-basin or Alignment (§2.2.3) Repermutate multiple models into the same basin
@@
Subspace-based Merging Methods (§2.3.3) Merge multiple models by projecting them into a sparse or low-rank subspace
Optimization-based Merging Methods (§2.3.4) Construct data-free optimization objectives to mitigate model conflicts and interference
@@
2.1. Notation and Model Merging Problem Definition
@@
In this survey, we focus on parameter-wise merging. In other words, the goal of model merging is to merge the parameters $\small\{\Theta^{(1)},\ldots,\Theta^{(T)}\}$, and finally obtain the new parameters $\small\Theta^{(merge)}=\texttt{merge}(\Theta^{(1)},\ldots,\Theta^{(T)})$. One straightforward solution for merging models is weighted averaging (Utans, 1996; Shoemake, 1985), defined as $\small\Theta^{(merge)}=\frac{1}{T}\sum_{t=1}^{T}\Theta^{(t)}$. However, the performance of this approach is often unacceptably poor or infeasible due to several possible factors: (i) The lack of suitable merging conditions, such as multiple models not being in the same basin or having inconsistent architectures. (ii) There are conflicts and interference among multiple models. We illustrate how advanced methods address these issues in §2.2 and §2.3, respectively.
2.2. Pre-Merging Methods
To better prepare models for merging, one class of work focuses on the fine-tuning and preprocessing of expert models, including linear fine-tuning to disentangle weights and sparse fine-tuning to reduce parameter conflicts during subsequent merging2.2.1), transforming heterogeneous architectures into a common one (§2.2.2), and rebasining or aligning weights before merging (§2.2.3).
2.2.1. Merger-Friendly Fine-Tuning
@@
Merge-friendly fine-tuning designs training strategies that produce models more amenable to merging than standard fine-tuning, as shown in Figure 3(a), with minimal performance degradation after merging. (i) Linearization Fine-tuning. Ortiz-Jimenez et al. (2023) reveal that one necessary condition for effective model merging is ‘weight disentanglement’. This means that different directions of the weight space correspond to functional changes in disjoint regions of the input space. To achieve weight disentanglement, Ortiz-Jimenez et al. (2023) propose fine-tuning the linearized model along the tangent space (Jacot et al., 2018) of the pre-trained model during the fine-tuning stage, rather than in the original space of the nonlinear model. However, linearized fine-tuning with all parameters is more expensive than nonlinear fine-tuning. To accelerate this process, some works suggest linearizing only part of the layers. For example, Tang et al. (2024c) propose partially linearizing the Adapter modules and then merging Adapters. Jin et al. (2024) suggest linearly fine-tuning only the linear layers in the attention modules of the full model. Furthermore, TAFT (Liu et al., 2024b) develops an efficient linearization method for the Transformer (Vaswani et al., 2017) architectures, which directly derives closed-form linearized solutions for transformer networks. (ii) Sharpness-Aware Fine-Tuning. SAFT-Merge (Lee et al., 2025) adopts a sharpness-aware minimization to encourage a flatter task-specific loss landscape, thereby reducing the sensitivity of model parameters to merging-induced perturbations and better preserving performance after model merging. (iii) Subspace Fine-Tuning. Unlike the full-parameter fine-tuning strategies mentioned above, some approaches fine-tune only a subset of parameters instead of the whole set. Keeping most weights frozen during fine-tuning can reduce conflicts that arise in subsequent merging. For example, OSRM (Zhang and Zhou, 2025) fine-tunes the model via LoRA in an orthogonal subspace, thereby mitigating unintended interference across tasks. In addition, TaLoS (Iurada et al., 2025) sparsely fine-tunes only a selected subset of parameters in the pretrained model to construct sparse task vectors with minimal interference.
@@
Figure 3. (a) Illustration of merge-friendly fine-tuning: unlike standard fine-tuning, which yields models that are hard to merge, merge-friendly fine-tuning produces models that are easier to merge. (b) Illustration of an architectural transformation that converts multiple heterogeneous models into a homogeneous architecture, enabling direct parameter-level merging. (c) Illustration of weight/parameter alignment: permuting the neural network model $\small\Theta^{(1)}$ so that it aligns with $\small\Theta^{(2)}$.
2.2.2. Architecture Transformation
In some cases, models that need to be merged may have different architectures and cannot be merged directly. To solve this problem, some studies (Avrahami et al., 2022; Wan et al., 2024a, b; Nguyen et al., 2023) propose to perform architecture transformation before merging, that is, transform multiple models with different architectures into the same architecture as shown in Figure 3(b). For example, GAN Cocktail (Avrahami et al., 2022) attempts to merge multiple GAN models $\{\Theta^{(1)},\ldots,\Theta^{(T)}\}$ with different architectures. It transforms all GAN models $\small\Theta^{(t)}$ ($t\in\{1,2,\ldots,T\}\;\&\;t\!\neq\!k$) into a specified target model $\Theta^{(k)}$, that is, $\small\Theta^{(k)}$ is used as the initialization to learn the output of $\small\Theta^{(t)}$, while adding implicit regularizations to ensure that $\small\Theta^{(k)}$ does not forget knowledge of the task $k$. Similarly, FuseLLM (Wan et al., [n. d.]) and FusionChat (Wan et al., 2024b) propose to merge chat LLMs with diverse architectures and scales (e.g., NH2-Mixtral-8x7B (Jiang et al., 2024), NH2-Solar-10.7B (Kim et al., 2023), OpenChat-3.5-7B (Wang et al., 2023a) in their practical applications). Specifically, FusionChat first uses knowledge distillation to transform all the architectures to match that of OpenChat-3.5-7B, and then performs the merge operation. Unlike the above distillation-based approach, CLAFusion (Nguyen et al., 2023) adds layers/blocks (with weights set to the identity matrix) to the smaller model to align its architecture with that of the larger model. In summary, merging models with different architectures requires first transforming the models into a common architecture.
2.2.3. Weight Re-basin or Alignment